OpenStudy (anonymous):

y varies inversely with both x and z. y=3 when x=4 and z=0.05

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

Write an equation that describes the following variation.

OpenStudy (anonymous):

z = -y/60 and x = -4y/3 => y = -3x/2 -30z

OpenStudy (anonymous):

its wrong :( but tnx for ur help

OpenStudy (anonymous):

So what's the answer then? If you know it's wrong, you must know the answer...

hartnn (hartnn):

xyz=k k=4*3*0.05=0.6 xyz=0.6

hartnn (hartnn):

y varies inversely with both x and z y= k/(xz) thats why xyz=k

OpenStudy (anonymous):

Nice thanks!